Ciaran is twice as old as Brid.30 years ago,Ciaran was 12 times as old as Brid.How old are they know
Answer by rothauserc(4718)

let C be Ciaran's age and B be Brid's age, then
:
1) C = 2 * B
:
2) C - 30 = 12 * (B - 30)
:
use equation 1 and substitute for C in equation 2
:
2B - 30 = 12B - 360
:
10B = 330
:
B = 33
C = 2 * 33 = 66
:
****************************
Ciaran is 66 and Brid is 33
****************************
: